OpenStudy (anonymous):

Give the y intercepts for the ellipse 16x^2+9y^2=144

OpenStudy (anonymous):

center : (0,0) division both side by 144 you get

OpenStudy (anonymous):

y intercept = + - 4

OpenStudy (anonymous):

\[\frac{x^2}{9}+\frac{y^2}{16}=1\]

OpenStudy (anonymous):

y=16 r=4 0+- 4 = +- 4

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

away wc

OpenStudy (anonymous):

Set x to zero in the equation. 9 y^2 = 144 y = 4 and y = -4 The y intercept are (0,4) and (0,-4) Refer to the attachment.